the ppc for country producing two good x and y giving by the formular y=100-20x assuming the is full employment from resourses, what is the opportunity cost of producing one extra unit of x.
1
Expert's answer
2014-10-03T15:04:22-0400
y=100-20x
The opportunity cost of a choice is the value of the best alternative forgone, in a situation in which a choice needs to be made between several mutually exclusive alternatives given limited resources.
The opportunity cost of producing one extra unit of x is: y' = (100 - 20x) = 20 units of y.